OpenStudy (anonymous):

Can someone check my answers? to the following 1) The Decomposition of potassium chlorate KCLO3 is used as a source of oxygen in the laboratory. How many moles of KCL3 are needed to produce 15 mol of oxygen. My answer: KCLO3----> KCL+O2 KCLO3----> 2KCL+3O2 15 mol of O* 2 mol Kclo3 per 3 mol= moles of KCLO3 moles of O* moles of KCLO3/ moles of O(15 moles of o*2 moles kclo3)/3 mol O= 10 mol KCLo3 2) when copper metal is added to silver nitrate in a solution silver metal and copper 2 nitrate are produced. mass of silver produced from 100. g of cu 2ag+cu-->cu2+2ag 339.5 g sliver

OpenStudy (anonymous):

3) If 20.5g chlorine is reacted with 20.5 g na wich reactant is excess ow do you know. cl2+2na-->2nacl so cl2=na1;2 cl2 mass 20.5/70.9=0.289 na mass 20.5/23= 0.891 sodium is exess cause there more of it

OpenStudy (anonymous):

3) If 20.5g chlorine is reacted with 20.5 g na wich reactant is excess ow do you know. cl2+2na-->2nacl so cl2=na1;2 cl2 mass 20.5/70.9=0.289 na mass 20.5/23= 0.891 sodium is exess cause there more of it

OpenStudy (anonymous):

Carbon reacts with steam h20 at high temptures to produce hydrogen and carbon monoxide if 2.40 mol carbon are exposed to 3.10 steam identify limiting My work- c+h20--> co+h2 2.4 moles carbon/1=2.4 water is 3.1/1 so carbon is limiting and water is excess

OpenStudy (anonymous):

@zepdrix @iheartfood

OpenStudy (anonymous):

@zoomiemoo